OpenStudy (solomonzelman):

Desperately need HELP !!! I posted this question before and what I got was "trig .... read algebra book!" or something along those lines!

OpenStudy (solomonzelman):

I want to know all the formulas/theorems that are needed to solve limits that involve trigonometric functions.

OpenStudy (solomonzelman):

Link, or type them in. ANYONE?

OpenStudy (solomonzelman):

I need rules like \[\huge\color{blue}{\lim_{x \rightarrow 0} ~~~ \frac{Tan(x) }{x} =1 } \] I need all, or as many as possible

OpenStudy (turingtest):

the only two I ever use are \[\lim_{x\to0}{\sin x\over x}=1\]and\[\lim_{x\to0}{1-\cos x\over x}=0\]

OpenStudy (turingtest):

all others that I have ever run across can be derived in terms of those two

OpenStudy (solomonzelman):

Good, ty! Appreciate your help - not more or less :)))))) !

OpenStudy (turingtest):

welcome :) hopefully you see that the limit in your example follows from the first I gave

OpenStudy (solomonzelman):

Alright, I don't have "my example" though, I am learning in advance sort of. Anyway, good luck!

OpenStudy (solomonzelman):

that's how I knew to derive and L'H'S :)

OpenStudy (turingtest):

i meant \[{\lim_{x \rightarrow 0} ~~~ \frac{\tan(x) }{x} =1 }\]

OpenStudy (solomonzelman):

yeah, I figured )) Thank you, and see you soon !